# Normalized mutual information is a biased measure for classification and community detection

### v2

@@ III Example application:
We perform community detection using six well-known algorithms as follows. (We use the implementations found in the igraph library [16], except for the inference method, for which we use the graph-tool library [17].)
@@
    Modularity maximization: Modularity is a quality function for network labelings equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving the node degrees. The labeling with the highest modularity is considered the best community division. Modularity can be maximized using a number of heuristic methods, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [19, 20], spectral methods [21], and simulated annealing [22, 23, 24]. In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
@@
As described in Section III.2, all of these algorithms perform reasonably well, but, as we will see, the best performers in the context of the current tests are InfoMap and the two variants of modularity maximization.
III.1 Comparison between variants of the
mutual information
Figure 4 summarizes the relative performance of the various mutual information measures in our tests. In this set of tests we limited ourselves, for the sake of clarity, to the top three community detection algorithms—InfoMap and the two variants of modularity maximization—and measured which of the three returned the best results according to each of our six mutual information measures, as a function of network size $n$ and the mixing parameter $\mu$. Each point in each of the six panels in the figure is color-coded with some mix of red, green, and blue to indicate in what fraction of cases each of the algorithms performs best according to each of the six measures and, as we can see, the results vary significantly among measures. An experimenter trying to choose the best algorithm would come to substantially different conclusions depending on which measure they use.
One consistent feature of all six mutual information measures is the large red area in each panel of Fig. 4, which represents the region in which InfoMap performs best. Regardless of the measure used, InfoMap is the best performer on networks with low mixing parameter (i.e., strong community structure) and relatively large network size. For higher mixing (weaker structure) or smaller network sizes, modularity maximization does better. Which version of modularity is best, however, depends strongly on the mutual information measure. The traditional symmetrically normalized mutual information (top left panel) mostly favors the version with a high resolution parameter of $\gamma=10$ (blue), but the asymmetric reduced measure for which we advocate (bottom right) favors the version with $\gamma=1$ (green). (The regions colored gray in the figure are those in which no algorithm receives a positive mutual information score and hence all algorithms can be interpreted as failing.)
These results raise significant doubts about the traditional measure. At large $n$ and high mixing $\mu$ the network will have many small communities, and in this regime it is natural that modularity maximization with a high resolution parameter will perform well because it favors large numbers of communities, and all mutual information measures agree on this point. For smaller system sizes or lower mixing, however, the network will have fewer, larger communities and in this regime we find that the $\gamma=10$ algorithm dramatically overestimates the number of communities in the network. With a non-reduced mutual information there is no penalty for doing this, as discussed in Section II.1, which is why the traditional normalized mutual information ranks the $\gamma=10$ algorithm highly in this regime, arguably erroneously. The reduced mutual information, by contrast, prefers the $\gamma=1$ algorithm. The number of groups inferred by each algorithm for each value of the mixing parameter $\mu$ is shown for $n=3200$ in Figure 5, which highlights these findings.
Figure 5: The number of groups inferred by each of the six algorithms in Fig. 4 for LFR benchmark networks with $n=3200$ nodes and a range of values of the mixing parameter $\mu$. The true number of groups is shown in black. The InfoMap algorithm (red) generates an accurate number of groups for values of $\mu$ up to about 0.5, but beyond this point it erroneously places all nodes in a single group. Standard modularity maximization with resolution parameter $\gamma=1$ (green) underestimates the number of groups, presumably because of the resolution limit on the detection of small groups, but not as severely as the number is overestimated when $\gamma=10$ (blue).
An extreme example of the disagreement between measures can be seen in the lower right corner of each plot in Fig. 4, where most measures return a negative value, indicating that community detection has failed altogether. This is expected: for very weak community structure all detection algorithms are expected to show a “detectability threshold” beyond which they are unable to identify any communities [30, 31, 32]. The standard normalized mutual information, however, claims to find community structure in this regime using the $\gamma=10$ version of modularity maximization. This occurs because the $\gamma=10$ algorithm finds many small communities and, as discussed in Section II.1, a labeling with many communities, even completely random ones, is accorded a high score by a non-reduced mutual information. This offers a clear reason to avoid the standard measure.
Figure 6: Performance of each of the six community detection algorithms considered here in identifying the known communities in a large set of LFR benchmark networks, as quantified by the asymmetrically normalized reduced information measure $I^{(A)}(c;g)$ proposed in this paper.
The bottom left panel in Fig. 4 shows results for the asymmetrically normalized version of the traditional mutual information, which gives even worse results than the symmetric version, with hardly any region in which the $\gamma=1$ version of modularity maximization outperforms the $\gamma=10$ version. This behavior arises for the reasons discussed in Section II.2: the bias inherent in the symmetric normalization fortuitously acts to partially correct the errors introduced by neglecting the information content of the contingency table. The asymmetric normalization eliminates this correction and hence performs more poorly. The correct solution to this problem, however, is not to use a symmetric normalization, which can bias outcomes in other ways as we have seen, but rather to adopt a reduced mutual information measure.
Finally, comparing the middle and right-hand columns of Fig. 4, we see that the results for the adjusted and reduced mutual information measures are quite similar in these tests, although there are some differences. In particular, the adjusted measure appears to find more significant structure for higher mixing than the reduced measure. This occurs because, as discussed in Section II.1 and Ref. [6], the adjusted measure encodes the contingency table in a way that is optimized for more uniform tables than the reduced measure, and thus penalizes uniform tables less severely, leading to overestimates of the mutual information in the regime where detection fails—in this regime the candidate and ground-truth labelings are uncorrelated which results in a uniform contingency table. This provides further evidence in favor of using a reduced mutual information measure.
III.2 Comparison between
community detection algorithms
Figure 6 shows results for all six algorithms listed in Section III. Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model. Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better. As noted by previous authors [33], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime. In this regime one must use another algorithm. Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore. (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
Thus our recommendations for the best community detection algorithm are relatively straightforward, if we base our choice on performance under the LFR benchmark, as is commonly done in this field. If we are in a regime where InfoMap succeeds, meaning it finds more than one community, then one should use InfoMap. If not, one should use standard modularity maximization with $\gamma=1$. That still leaves open the question of how the modularity should be maximized. In our studies we find the best results with simulated annealing, but simulated annealing is computationally expensive. In regimes where it is not feasible, we recommend using the Leiden algorithm instead. (Tests using other computationally efficient maximization schemes, such as the Louvain and spectral algorithms, generally performed less well than the Leiden algorithm.)
IV Conclusions
In this paper we have examined the performance of a range of mutual information measures for comparing labelings of objects in classification, clustering, or community detection applications. We argue that the commonly used normalized mutual information is biased in two ways: (1) because it ignores the information content of the contingency table, which can be large, and (2) because the symmetric normalization it employs introduces spurious dependence on the labeling. We argue in favor of a different measure, an asymmetrically normalized version of the so-called reduced mutual information, which rectifies both of these shortcomings.
Acknowledgements.
The authors thank Samin Aref for useful comments and feedback. This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617, and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
Appendix A Benchmarking
In this appendix we give some additional details of our numerical tests.
A.1 Results for the traditional symmetric normalized mutual information
@@
Figure 7: Performance of each of the six community detection algorithms considered here, as quantified by the conventional symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$.
@@
Figure 7 shows the performance of the same six community detection methods as in Fig. 6, but measured using the standard, symmetrically normalized, non-reduced mutual information $I_{0}^{(S)}(c;g)$. By this measure many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig. 6 in the regime where all methods should by rights fail. Note in particular the high scores achieved by the generalized modularity with $\gamma=10$ in this undetectable regime by virtue of the excessive number of groups it generates.
A.2 LFR network generation
@@
The LFR networks we use for benchmarking are generated using the procedure described in [15], which is as follows.
@@
    Fix the number of nodes $n$ and mixing parameter $\mu$. In our studies we use node counts in the range $n\in[200,51200]$. The parameter $\mu$ controls the relative number of edges within and between communities. For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect. But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [30, 34]. In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
    Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$. Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 [36, 37, 38, 15]. We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$. Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
@@
    Assign each node to a community randomly, one node at a time, while ensuring that the community chosen is always large enough to support the added node’s intra-community degree, given by $(1-\mu)k$ where $k$ is the total degree.
@@
The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) [26], and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks. Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.). By using the LFR benchmark with parameters $\theta$ to compare the performance of community detection algorithms, we are therefore implicitly defining the “best” algorithm to be the one whose corresponding function $h[A]$ optimizes $\sum_{A,g}P(A,g|\theta)M(g,h[A])$. If the similarity measure we choose is the “all or nothing” error function $M(g,c)=\delta(g,c)$, then the optimal community detection algorithm is trivially the one with
@@
In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated. There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results [39]. The LFR model, however, is not precisely a special case of the DCSBM. In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law. Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion. These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.

### v3

@@ A.2 Community detection algorithms
We assess the performance of six common community detection algorithms on synthetic network examples. We use the algorithm implementations from the igraph library [26], except for the inference method, for which we use the graph-tool library [27]. The six algorithms are as follows.
@@
    Modularity maximization: Modularity is a quality function for network community divisions equal to the fraction of edges within communities minus the expected such fraction if edge positions are randomized while preserving node degrees. The labeling with the highest modularity is considered the best community division. Exact modularity maximization is NP-hard and usually intractable, but modularity can be maximized approximately using various heuristics, of which the most popular are agglomerative methods such as the Louvain and Leiden algorithms [29, 30], spectral methods [31], and simulated annealing [32, 33, 34]. In our tests we use simulated annealing where computationally feasible and the Leiden algorithm otherwise, these approaches giving the most consistent maximization of the modularity.
@@
As discussed in Section 2.5, all of these algorithms perform reasonably well, but the best performers in our tests are InfoMap and the two variants of modularity maximization.
A.3 Results for the traditional symmetric normalized mutual information
@@
Figure 8: Performance of each of the six community detection algorithms considered here, as quantified by the conventional symmetrically normalized non-reduced mutual information $I_{0}^{(S)}(c;g)$.
@@
Figure 8 shows the performance of the same six community detection methods as in Fig. 6, but measured using the standard, symmetrically normalized, non-reduced mutual information, which we denote by $I_{0}^{(S)}(c;g)$. By this measure, many of the methods appear to perform implausibly well, far beyond the detectability threshold visible in Fig. 6, in the regime where all methods should by rights fail. Note in particular the high scores achieved by the generalized modularity with $\gamma=10$ by virtue of the excessive number of groups it generates.
A.4 LFR network generation
@@
The LFR networks we use for benchmarking are generated using the procedure described in [20], which we summarize here.
@@
    Fix the number of nodes $n$ and mixing parameter $\mu$. In our studies we use node counts in the range $n\in[200,51200]$. The parameter $\mu$ controls the relative number of edges within and between communities. For small $\mu$ there are many more edges within communities than between them, which makes the communities easy to detect. But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [21, 40], although there is no known analytic detectability threshold for the LFR model. In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
    Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$. Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 [42, 43, 44, 20]. We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left(n/10,100\right)$. Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
@@
    Assign each node to a community randomly, one node at a time, ensuring that the community chosen is large enough to support the added node’s intra-community degree, given by $(1-\mu)k$ where $k$ is the total degree.
@@
The LFR model is similar to a special case of the degree-corrected stochastic block model (DC-SBM) [36], and hence one might expect that inference-based community detection methods employing the latter model would perform well, perhaps even optimally, on LFR networks. Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.). By using the LFR benchmark with parameters $\theta$ to compare the performance of community detection algorithms, we are therefore implicitly defining the “best” algorithm to be the one whose corresponding function $h[A]$ optimizes $\sum_{A,g}P(A,g|\theta)M(g,h[A])$. If the similarity measure we choose is the “all or nothing” error function $M(g,c)=\delta(g,c)$, then the optimal community detection algorithm is trivially the one with
@@
In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated. There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DC-SBM, we might expect the DC-SBM (with appropriate priors) to give optimal results [45].
The LFR model, however, is not precisely a special case of the DC-SBM. In particular, the DC-SBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law. Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregate and weigh different modes of error in a sensible fashion. These differences, it appears, are enough to ensure that the DC-SBM does not perform the best in our testing.
@@
Appendix B Bias towards large numbers of groups in the traditional mutual information
To shed light on why the traditional mutual information of Eq. (5) is biased towards an excessive number of groups and how the reduced mutual information of Eq. (4) corrects this, consider the (extreme) example in which every object is placed in a group on its own, producing a candidate labeling $c=(1,\ldots,n)$. Whatever the ground truth labeling $g$ is, this candidate labeling clearly provides no information about it whatsoever, so we expect the mutual information to be zero. But this is not what we find. The contingency table in this case is
[…]
so the conventional mutual information of Eq. (5) is
$I_{0}(c;g)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}=H_{0}(g).$ (16)
This answer is not merely wrong; it is maximally so. The mutual information should take its minimum value of zero, but instead it takes the value $H_{0}(g)$, which is the maximum possible since $H_{0}(g)$ is an upper bound (see Eq. (2)). The reason for this result is that in this case the contingency table itself uniquely defines $g$, so neglecting it puts the mutual information in error by an amount equal to the complete information cost of the ground truth. If we include the the information cost of the table, on the other hand, as we do in the reduced mutual information, this erroneous behavior disappears. Assuming for simplicity that the ground-truth groups are of equal size, the optimal concentration parameters are $\alpha_{g|c}=0$ and $\alpha_{g}=\infty$, and the reduced mutual information of Eq. (4) becomes
$=0.$ (17)
which is the correct answer.
Outside this special case, we observe from Eq. (4) that in the limit where the numbers of groups $q_{g}=q_{c}=q$ are held fixed and $n\rightarrow\infty$, the traditional mutual information grows as
$I_{0}(c;g)\sim\mathrm{O}(n\log q)$ (18)
while the correction term grows as
$I(c;g)-I_{0}(c;g)\sim\mathrm{O}(q^{2}\log n).$ (19)
Thus, in this limit the subleading term is relevant when $n/\log n\lesssim q^{2}/\log q$, a condition that holds in many practical contexts, where the number of communities can grow as $\sqrt{n}$ or faster.